# POGEMA: A Benchmark Platform for Cooperative Multi-Agent Pathfinding

### v2

@@ 4.3 Experimental Results
For MAPF tasks, LaCAM notably outperforms all other approaches. Specialized learnable MAPF approaches, such as DCC and SCRIMP, follow closely, with similar overall performance but different strengths. DCC performs better than SCRIMP on out-of-distribution tasks and pathfinding tasks, while the rest of the results are fairly similar. Surprisingly, SCRIMP underperforms on pathfinding tasks, indicating a weakness in single-agent tasks that do not require communication, which may be an out-of-distribution scenario for this algorithm. It is also worth noting that SCRIMP includes an integrated tie-breaking mechanism that ensures collision-free actions. MARL algorithms such as QPLEX, VDN, and QMIX perform significantly worse than other approaches, showing a substantial gap in the results. This can be attributed to the absence of additional techniques used in hybrid approaches. This may suggest that the MARL community lacks large-scale approaches and benchmarks for these tasks. Among MARL approaches, MAMBA achieves the best results in performance, cooperation, and pathfinding metrics, which can be attributed to its communication mechanism. However, its performance remains much worse than that of specialized methods, and it fails to solve any instances in the out-of-distribution dataset.
@@ B.1 Performance
The performance metrics were calculated using Mazes and Random maps of size close to $20\times 20$. The primary metrics here are SoC and CSR. The results of all the MAPF approaches over different numbers of agents are presented in Figure 4. The superior performance is shown by the centralized approach, LaCAM. The learnable approaches, DCC and SCRIMP, show comparable results. Interestingly, the former has a better SoC metric, despite the latter having better results on CSR. Among the MARL methods, better results are shown by MAMBA for both metrics. However, it narrowly lags behind the specialized approaches, DCC and SCRIMP.
@@ B.2 Out-of-Distribution
Out-of-Distribution metric was calculated on Cities-tiles dataset, that consists of pieces of cities maps with $64\times 64$ size. Due to much larger size compared to Mazes and Random maps, the amount of agents was also significantly increased. The results are presented in Figure 5. Here again centralized search-based planner, i.e. LaCAM, demonstrates the best results both in terms of CSR and SoC. Among hybrid methods, DCC completely outperforms the others (including SCRIMP) on the out-of-distribution dataset. MARL approaches are unable to solve any instance even with 64 agents.
@@ B.3 Scalability
The results of how well the algorithm scales with a large number of agents are shown in Figure 6. The experiments were conducted on a warehouse map. The plot is log-scaled. The best scalability is achieved with the centralized LaCAM approach, which is a high-performance approach. The worst results in both runtime and scalability are for SCRIMP, with results close to it for DCC. Despite an initially high runtime, the scalability of MAMBA is better than other approaches; however, this could be attributed to the high cost of GPU computation, which is due to the large number of parameters in the neural network and is the limiting factor of this approach.
@@ B.4 Cooperation
DCC $0.75$ p m 0.04 $89.36$ p m 10.34
IQL $0.47$ p m 0.02 $250.03$ p m 18.83
LaCAM $1.00$ p m 0.00 $20.84$ p m 1.75
MAMBA $0.40$ p m 0.02 $173.00$ p m 13.92
QMIX $0.35$ p m 0.04 $253.28$ p m 16.24
QPLEX $0.51$ p m 0.05 $234.79$ p m 18.89
SCRIMP $0.85$ p m 0.03 $93.18$ p m 13.23
VDN $0.45$ p m 0.05 $242.42$ p m 16.87
@@
How well the algorithm is able to resolve complex situations on the Puzzles set is reflected in the results presented in Table 4. Surprisingly, the centralized approach LaCAM does not solve all the tasks, showing only a $0.96$ CSR. This highlights that this type of task is difficult even for centralized approaches, despite the small map size of $5\times 5$ and the low number of agents ($2-4$). SCRIMP outperformed DCC in CSR but again showed comparable results in SoC. Among MARL approaches, better cooperation is demonstrated by QMIX, outperforming QPLEX, VDN, IQL, and even MAMBA.
@@ Appendix I Resources and Statistics
For this test, we used default observation parameters commonly employed in learnable multi-agent pathfinding (MAPF) approaches (e.g., Follower, SCRIMP, DCC). Specifically, we set the observation radius to 5, corresponding to an $11\times 1$ 11 field. The results for both MAPF and LMAPF scenarios, using a random policy, are reported in Table 12. In scenarios with more than 32 agents, POGEMA achieves $\geq 80$K observations per second (OPS) on fast CPUs like the Ryzen Threadripper and Apple Silicon M1, and $\geq 38$K OPS on the Intel Xeon setup, which is notably fast for a single-environment, single-thread configuration. For comparison, EnvPool [^20] https://github.com/sail-sg/envpool reports 50K frames per second (FPS) for Atari on a 12-CPU setup.

### v3

@@ 4.3 Experimental Results
For MAPF tasks, LaCAM notably outperforms all other approaches. Specialized learnable MAPF approaches, such as DCC and SCRIMP, perform better than any of the MARL approaches. When comparing these two specialized approaches, SCRIMP outperforms DCC across such metrics as Performance, Cooperation, and Out-of-Distribution. Surprisingly, SCRIMP underperforms on pathfinding tasks, revealing a weakness in single-agent scenarios that don’t require communication, which may represent an out-of-distribution case for this algorithm. It is also worth noting that SCRIMP incorporates an integrated tie-breaking mechanism that ensures collision-free actions. MARL algorithms such as QPLEX, VDN, and QMIX perform significantly worse than other approaches, showing a substantial gap in the results. This can be attributed to the absence of additional techniques used in hybrid approaches. This may suggest that the MARL community lacks large-scale approaches and benchmarks for these tasks. Among MARL approaches, MAMBA achieves the best results in performance, cooperation, and pathfinding metrics, which can be attributed to its communication mechanism. However, its performance remains much worse than that of specialized methods, and it fails to solve any instances in the out-of-distribution dataset.
@@ Appendix B Results for MAPF Benchmark
Please note that the previous arXiv version contains different SCRIMP results due to a technical error, which we discovered while refactoring the code for the camera-ready version.
@@ B.1 Performance
The performance metrics were calculated using Mazes and Random maps of size close to $20\times 20$. The primary metrics here are SoC and CSR. The results of all the MAPF approaches over different numbers of agents are presented in Figure 4. The superior performance is shown by the centralized approach, LaCAM. Next best-performing approach is SCRIMP that is substantially outperforms another specialized solver – DCC. Among the MARL methods, better results are shown by MAMBA for both metrics. However, it narrowly lags behind the specialized approaches, DCC and SCRIMP.
@@ B.2 Out-of-Distribution
Out-of-Distribution metric was calculated on Cities-tiles dataset, that consists of pieces of cities maps with $64\times 64$ size. Due to much larger size compared to Mazes and Random maps, the amount of agents was also significantly increased. The results are presented in Figure 5. Here again centralized search-based planner, i.e. LaCAM, demonstrates the best results both in terms of CSR and SoC. Among hybrid methods, SCRIMP demonstrates better performance in terms of CSR, however, average SoC of its solutions is almost the same as solutions found by DCC. MARL approaches are unable to solve any instance even with 64 agents.
@@ B.3 Scalability
The results of how well the algorithm scales with a large number of agents are shown in Figure 6. The experiments were conducted on a warehouse map. The plot is log-scaled. The best scalability is achieved with the centralized LaCAM approach, which is a high-performance approach. DCC demonstrates the worst results in terms of runtime; however, regarding the Scalability metric, DCC’s results are identical to SCRIMP’s. Despite an initially high runtime, the scalability of MAMBA is better than other approaches; however, this could be attributed to the high cost of GPU computation, which is due to the large number of parameters in the neural network and is the limiting factor of this approach.
@@ B.4 Cooperation
DCC $0.74$ p m 0.04 $93.50$ p m 10.47
IQL $0.47$ p m 0.04 $250.03$ p m 19.09
LaCAM $1.00$ p m 0.00 $20.84$ p m 1.65
MAMBA $0.40$ p m 0.05 $173.60$ p m 14.04
QMIX $0.35$ p m 0.04 $253.28$ p m 16.04
QPLEX $0.51$ p m 0.05 $234.79$ p m 19.15
SCRIMP $0.85$ p m 0.03 $77.86$ p m 11.09
VDN $0.45$ p m 0.05 $242.42$ p m 17.92
@@
How well the algorithm is able to resolve complex situations on the Puzzles set is reflected in the results presented in Table 4. The only approach out of the evaluated ones that is able to solve all the instances in this set is LaCAM. Out of the rest approaches, as well as on other sets of instances, the best results is demonstrated by SCRIMP, which outperforms DCC both in terms of CSR and SoC. Among MARL approaches, better cooperation is demonstrated by QMIX, outperforming QPLEX, VDN, IQL, and even MAMBA.
@@ Appendix I Resources and Statistics
For this test, we used default observation parameters commonly employed in learnable multi-agent pathfinding (MAPF) approaches (e.g., Follower, SCRIMP, DCC). Specifically, we set the observation radius to 5, corresponding to an $11\times 1$ 11 field. The results for both MAPF and LMAPF scenarios, using a random policy, are reported in Table 12. In scenarios with more than 32 agents, POGEMA achieves $\geq 80$K steps per second (OPS) on fast CPUs like the Ryzen Threadripper and Apple Silicon M1, and $\geq 38$K OPS on the Intel Xeon setup, which is notably fast for a single-environment, single-thread configuration. For comparison, EnvPool [^20] https://github.com/sail-sg/envpool reports 50K frames per second (FPS) for Atari on a 12-CPU setup.